# A parameterized block-splitting preconditioner for indefinite least squares problem

### v1

@@ 4 Numerical experiments
where the function $g$ is known.The discretization of this equation employing the second-order central difference scheme on an $(n_{0}+2)$ grid in all directions of $\Omega$, with a mesh size of $h=1/(n_{0}+1)$, yields a system of linear equations with the coefficient matrix $A_{1}\in\mathbb{R}^{n\times n},$ in which $n=n_{0}^{2}$. We set $A_{2}=0.7I_{n}$, where $I_{n}$ is the identity matrix of order $n$.
We set $n_{0}=85,90$ and $95$. Consequently, the resulting matrices $A_{1}$ are of orders $7225$, $8100$, and $9025$, respectively. In this case, the order of matrices $A$ is $14450\times 7225$, $16200\times 8100$ and $18050\times 9025$. All the assumptions are as the previous example, except here we have used the complete version of GMRES. All the other assumptions are the same as in the previous example. Numerical results are presented in Table 3. As we observe the PBS preconditioner outperforms the other preconditioners.
@@
   $m\times n$ PBS BS1 BS2 BS3 No-Prec
[…]
CPU 0.10 0.18 0.11 0.20
RES 7.49e-12 5.86e-12 4.63e-13 1.47e-12
ERR 4.30e-09 1.01e-09 1.68e-09 1.68e-09
[…]
CPU 0.11 0.21 0.13 0.23
RES 7.52e-12 5.87e-12 5.22e-13 1.65e-12
ERR 3.43e-09 1.88e-09 1.88e-09 1.88e-09
[…]
CPU 0.13 0.24 0.15 0.27
RES 7.53e-12 5.88e-12 5.71e-13 1.75e-12
ERR 5.85e-09 1.65e-09 1.64e-09 1.65e-9
@@ 5 Conclusions
We have presented a parameterized block splitting iteration method for solving indefinite least squares problem. We have proved that the iteration method is convergent under a condition. Optimal value of the parameter has been computed. Eigenpair properties of the preconditioned matrix has been studied. We have compared the numerical results of the preconditioner induced by the proposed method with three recently presented preconditioners. The numerical results show that the proposed preconditioner in more efficient that others.

### v3

@@ 4 Numerical experiments
where the function $g$ is known.The discretization of this equation employing the second-order central difference scheme on an $(n_{0}+2)$ grid in all directions of $\Omega$, with a mesh size of $h=1/(n_{0}+1)$, yields a system of linear equations with the coefficient matrix $A_{1}\in\mathbb{R}^{n\times n},$ in which $n=n_{0}^{2}$. We set $A_{2}=0.7I_{n}$, where $I_{n}$ is the identity matrix of order $n$. We set $n_{0}=85,90,95,100,105$ and $110$. Consequently, the resulting matrices $A_{1}$ are of orders $7225$, $8100$, $9025$, $10000$, $11025$ and $12100$, respectively. In this case, the order of matrices $A$ is $14450\times 7225$, $16200\times 8100$, $18050\times 9025$, $0000\times 10000, 22$$2050\times 11025$ and $24200\times 12100$. All the assumptions are as the previous example, except here we have used the complete version of GMRES. All the other assumptions are the same as in the previous example. Numerical results are presented in Table 3. As we observe the PBS preconditioner outperforms the other preconditioners.
@@
$n_{0}$    $m\times n$ PBS BS1 BS2 BS3 BS4 No-Prec
85 $14450\times 7225$ Iter 4 9 6 9 5 $\dagger$}
CPU 0.49 0.81 0.59 0.74 0.53
RES 7.49e-12 5.86e-12 4.63e-13 1.47e-12 1.43e-12
ERR 4.30e-09 1.01e-09 1.68e-09 1.68e-09 1.64e-9
90 $16200\times 8100$ Iter 4 9 6 9 5 $\dagger$}
CPU 0.57 0.96 0.69 0.86 0.62
RES 7.52e-12 5.87e-12 5.22e-13 1.65e-12 1.56e-12
ERR 3.43e-09 1.88e-09 1.88e-09 1.88e-09 1.28e-9
95 $18050\times 9025$ Iter 4 9 6 9 5 $\dagger$}
CPU 0.67 1.10 0.80 1.01 0.72
RES 7.53e-12 5.88e-12 5.71e-13 1.75e-12 1.75e-12
ERR 5.85e-09 1.65e-09 1.64e-09 1.65e-9 2.16e-9
100 $20000\times 10000$ Iter 4 9 6 9 5 $\dagger$}
CPU 0.78 1.28 0.94 1.18 0.86
RES 7.55e-12 5.91e-12 6.67e-13 1.90e-12 1.94e-12
ERR 4.87e-09 2.29e-09 2.29e-09 2.29e-9 2.29e-9
105 $22050\times 11025$ Iter 4 9 6 9 5 $\dagger$}
CPU 0.90 1.46 1.09 1.39 0.98
RES 7.56e-12 5.91e-12 7.28e-13 2.09e-12 2.14e-12
ERR 4.69e-09 2.08e-09 2.08e-09 2.08e-9 2.08e-9
110 $24200\times 12100$ Iter 4 9 6 9 5 $\dagger$}
CPU 1.03 1.66 1.25 1.58 1.12
RES 7.57e-12 5.93e-12 8.09e-13 2.28e-12 2.31e-12
ERR 7.69e-09 3.19e-09 3.19e-09 3.19e-9 3.19e-9
@@ 5 Conclusions
We have presented a parameterized block splitting iteration method for solving indefinite least squares problem. We have proved that the iteration method is conditionally convergent. Optimal value of the parameter has been computed. Eigenpair properties of the preconditioned matrix has been studied. We have compared the numerical results of the preconditioner induced by the proposed method with three recently presented preconditioners. The numerical results show that the proposed preconditioner in more efficient that others.
Acknowledgements
The author sincerely thanks the reviewers for their meticulous review of the paper and for their constructive comments and suggestions